### v1

# UnLoc: Leveraging Depth Uncertainties for Floorplan Localization

Abstract
We propose UnLoc, an efficient data-driven solution for sequential camera localization within floorplans. Floorplan data is readily available, long-term persistent, and robust to changes in visual appearance. We address key limitations of recent methods, such as the lack of uncertainty modeling in depth predictions and the necessity for custom depth networks trained for each environment. We introduce a novel probabilistic model that incorporates uncertainty estimation, modeling depth predictions as explicit probability distributions. By leveraging off-the-shelf pre-trained monocular depth models, we eliminate the need to rely on per-environment-trained depth networks, enhancing generalization to unseen spaces. We evaluate UnLoc on large-scale synthetic and real-world datasets, demonstrating significant improvements over existing methods in terms of accuracy and robustness. Notably, we achieve $2.7$ times higher localization recall on long sequences (100 frames) and $16.7$ times higher on short ones (15 frames) than the state of the art on the challenging LaMAR HGE dataset.
Figure 1: The proposed UnLoc processes an input image sequence to predict the floorplan depth (in meters) and associated uncertainty for each image column. Using these predictions, it generates a probability distribution over potential $\text{SE}(2)$ camera poses and outputs the most likely one (blue arrow). The ground truth pose of the last frame is also shown (red arrow), overlapped by the predicted pose.
1 Introduction
Camera localization within indoor environments is a fundamental problem in computer vision, essential for applications in augmented reality and robotics. Accurate localization enables devices to understand their spatial context, facilitating tasks such as navigation, object interaction, and autonomous exploration. Traditional localization methods often rely on pre-built 3D models [41, 43, 30, 39, 36] or extensive image databases [44, 2, 23, 48], which are storage-intensive and require substantial maintenance, limiting their scalability to new or dynamically changing environments.
Floorplans offer a lightweight and readily available alternative for indoor localization. As 2D representations of spaces, floorplans are easy to obtain [29, 56] and remain unaffected by changes in appearance, such as furniture rearrangements or lighting variations. Recent methods have leveraged floorplans for localization by aligning images with the map [18, 32, 10], enabling devices to localize within new environments as long as a floorplan is available.
Among these, F3Loc [10] has emerged as a recent promising approach for sequential visual floorplan localization, significantly outperforming all previous baselines. It integrates observations over time using a histogram filter, achieving impressive accuracy and computational efficiency on a range of datasets. However, as promising as it is, F3Loc has notable limitations that hinder its practical deployment, which we aim to improve upon in this paper:
Lack of Uncertainty Modeling: F3Loc combines monocular and multi-view depth estimation. However, it assumes that the resulting predictions are all of similar accuracy, and it has no means to represent and predict the uncertainties in the depth predictions. In indoor environments, depth estimation is often unreliable in regions with glass walls, open doorways, or large, featureless walls. When fusing a sequence of predictions, not accounting for uncertainty, inaccurate depth predictions adversely affect the localization process, leading to erroneous pose estimates.
Dataset-Specific Depth Networks: F3Loc relies on a custom depth prediction network that must be trained separately for each dataset or environment. This per-dataset training requirement poses significant challenges for scalability and robustness. Collecting sufficient depth data for retraining the depth network for every new environment is impractical, especially when rapid deployment is desired.
This paper addresses these shortcomings by introducing a novel visual floorplan localization method that incorporates uncertainty estimation into monocular depth prediction and uses it to robustly fuse predictions from a sequence of images. Also, our method enables us to leverage pre-trained monocular depth models. Our contributions are:
  • Uncertainty-Aware Depth Prediction: We model floorplan depth predictions as explicit probability distributions, enabling us to represent the uncertainty associated with each estimate. This uncertainty-aware approach improves localization by weighting predictions according to their confidence levels, which is particularly beneficial in challenging areas prone to depth estimation errors.
  • Leveraging Off-the-Shelf Depth Models: Instead of training custom depth networks for each environment, we utilize state-of-the-art monocular depth networks pre-trained on large-scale datasets [53]. These models provide accurate depth predictions out-of-the-box, eliminating the need for extensive per-environment training and enhancing the method’s generalization to new settings.
The proposed UnLoc (see Fig. 1) demonstrates significant improvements in accuracy and robustness on multiple datasets compared to the state of the art.
2 Related work
Figure 2: Method overview. At timestep $t$, UnLoc aligns an image with gravity and processes it through a pretrained monodepth encoder. Features extracted at this step, along with a binary mask from the gravity alignment, are used to predict the floorplan depth $\mathbf{\hat{d}}_{t}$ and its uncertainty $\mathbf{\hat{b}}_{t}$ via a masked attention mechanism. These predictions are used to form equiangular rays to perform uncertainty-aware matching with the floorplan’s occupancy map. Finally, a histogram filter fuses the observation likelihood with the integrated past belief.
Visual localization is a fundamental problem in computer vision, addressed through various approaches. Traditional methods include image retrieval techniques [12, 21, 2, 23, 48], which find the most similar images in a database and estimate the pose of the query image based on the retrieved ones. Structure-from-Motion-based approaches [1, 44, 42, 36] build a 3D model of the environment and establish 2D-3D correspondences by matching local descriptors, computing camera poses using minimal solvers [27] and RANSAC [15] or its recent variants [4, 3]. Scene coordinate regression methods [8, 9] learn to regress the 3D coordinates of image pixels, while pose regression techniques [25, 24] use networks to predict a 6-DoF camera pose from input images directly. These methods often rely on pre-built 3D models that are storage-intensive and scene-specific, limiting their applicability in unseen environments.
To overcome this, floorplan-based localization methods have emerged, utilizing overhead images or floorplans to estimate the $\text{SE}(2)$ pose of the camera [50, 46]. These approaches can localize images in new scenes as long as a floorplan is provided. Floorplan localization is frequently associated with LiDAR sensors [33, 54, 58], which are impractical for widespread mobile device use. Alternative methods reconstruct 3D geometry using depth cameras [49] or visual odometry [34]. Some approaches extract geometric features like room edges to align with the floorplan [7, 28]. However, these methods often assume known camera or room height, which is not always feasible.
Recent learning-based methods aim to use only RGB images for floorplan localization. LaLaLoc [19] estimates the position of panoramic images in a floorplan by embedding map and image features into a shared space. LaLaLoc++ [18] removes the known camera and ceiling height assumption by directly embedding the floorplan. LASER [32] represents the floorplan as a set of points and uses PointNet [37] to embed the visible points for each pose, aligning them with image features in a shared space. PF-Net [22] integrates localization within a differentiable particle filtering framework, using a learned similarity between images and corresponding map patches. F3Loc [10] utilizes metric monocular depth prediction but requires training custom depth networks, which can limit generalizability and assumes that all predictions are of the same quality.
Sequence-based localization methods enhance robustness by integrating information over time. Bayesian filtering [13, 11, 22, 7, 31] is commonly used to fuse sequential observations, like particle [13] and histogram filters [45]. PF-Net [22] employs a differentiable particle filter for localization but relies on learned observation models that may not generalize well. F3Loc [10] uses sequential observations and integrates them employing a histogram filter.
Depth estimation provides valuable geometric information for localization [10]. Nowadays, there are many strong pre-trained monodepth models [38, 5, 16, 55, 20, 52, 53, 6], trained on large-scale datasets, which can provide accurate metric and/or relative depth predictions out-of-the-box without the need for per-scene training. In this work, we leverage off-the-shelf monodepth networks with uncertainty estimation to enhance image-based floorplan localization. By utilizing pre-trained monodepth networks, we avoid custom depth model training, improving generalization. Incorporating uncertainty estimation enables us to guide the localization process using the confidence of depth predictions, focusing on reliable areas and improving accuracy.
3 Floorplan Localization with UnLoc
3.1. Method Overview. We estimate the 2D pose $\mathbf{s}_{t}=[s_{x,t},s_{y,t},s_{\phi,t}]$ within a known 2D floorplan, where $s_{x,t}$ and $s_{y,t}$ denote position coordinates and $s_{\phi,t}$ represents orientation. Our approach takes as input a sequence of $t$ RGB images, relative poses between these images, gravity direction, camera intrinsics, and the geometric layout of the floorplan. To ensure out-of-the-box usability of the method, the input floorplans are provided solely as occupancy grid data without any semantic labels.
An overview of our method is in Fig. 2. Each image is aligned with gravity and processed by a pre-trained encoder [53] to extract features. These features, along with a binary mask from the gravity alignment, are input to a masked attention mechanism that predicts the floorplan depth, defined as the depth to the nearest occupied area in the floorplan. The predicted depth is then used to construct equiangular rays, which are matched against the floorplan to produce a localization estimate. Finally, we combine this estimate with the previous posterior localization using a histogram filter to compute the updated posterior estimate.
Our approach differs from prior methods in two main ways: (1) we model floorplan depth as an explicit probability distribution to quantify uncertainty, enhancing sequential localization when fusing predictions; (2) we use an off-the-shelf pre-trained monodepth network instead of a custom network that requires per-dataset training, thereby effectively leveraging models trained on extensive data without additional training efforts.
3.2. Gravity Alignment. Real-world applications often involve images captured from hand-held or head-mounted devices, resulting in arbitrary camera orientations. To address this, we preprocess images to align them with the gravity.
We achieve gravity alignment by utilizing the camera’s roll ($\psi$) and pitch ($\theta$) angles to define rotation matrices. The rotation matrix for roll is given as $\mathbf{R}_{x}(\psi)$ and for pitch as $\mathbf{R}_{y}(\theta)$. The combined rotation matrix from the gravity-aligned frame to the camera frame is calculated as $\mathbf{R}_{cg}=\mathbf{R}_{y}(\theta)\cdot\mathbf{R}_{x}(\psi)$. The inverse transformation, from the camera frame to the gravity-aligned frame, is given by: $\mathbf{R}_{gc}=\mathbf{R}_{cg}^{\top}$. Using these rotations, we compute a homography $\mathbf{H}$ to warp the image into the gravity-aligned frame $\mathbf{H}=\mathbf{K}\cdot\mathbf{R}_{gc}\cdot\mathbf{K}^{-1}$, where $\mathbf{K}$ is the camera intrinsic matrix.
This process also produces a binary mask indicating pixels that are invalid after alignment due to the warping. The gravity-aligned RGB image and the mask are then used for subsequent feature extraction.
Note that the gravity direction and camera intrinsics are usually accessible from sensors on smartphones and head-mounted devices. If this information is not directly available, methods such as GeoCalib [47] can be used to estimate both the camera intrinsics and gravity direction.
3.3. Feature Extraction. Recent visual floorplan localization methods [19, 18, 32, 10] rely on encoders pre-trained on ImageNet [14] for image classification tasks, such as ResNet-50 [17]. However, we posit that encoders trained on tasks more closely related to floorplan depth estimation could yield better performance. To explore this, we utilize encoders pre-trained on dense monocular depth estimation tasks. These encoders, optimized on large-scale depth datasets, provide superior features for floorplan depth estimation without requiring additional depth training from scratch. From the state-of-the-art models  [5, 16, 55, 20, 52, 53, 6], we select the recent Depth Anything V2 model [53] due to its state-of-the-art performance in both relative and metric depth estimation and low inference times compared to similarly accurate models [6]. Specifically, we use the encoder fine-tuned for indoor environments. We extract features from its last layer and apply bilinear interpolation to match the spatial dimensions the subsequent masked attention mechanism requires.
Let us note that the proposed pipeline is agnostic to the monodepth network, which could easily be replaced as better methods become available. We provide an ablation study on selected encoders in Sec. 4.2.
3.4. Masked Attention. We implement a masked attention mechanism inspired by [10] to predict floorplan depth, using the interpolated features and gravity alignment mask as inputs. Our model predicts two 1D vectors: $\mathbf{\hat{d}}_{t}$, representing floorplan depth estimates, and $\mathbf{\hat{b}}_{t}$, denoting the uncertainty associated with each estimate. The uncertainty quantification allows the model to account for varying confidence levels across regions in the image, particularly in challenging areas with ambiguous visual cues or distant objects.
Before inputting to the attention mechanism, we reduce the channel dimensions of the interpolated encoder features using a convolutional layer. The resulting features serve as keys and values in the attention mechanism, while 1D queries are formed through average pooling. Positional encodings for the queries are derived from their 1D coordinates, whereas for the keys and values, positional encodings are mapped from the corresponding 2D image coordinates. By applying the gravity alignment mask, we focus the attention mechanism on observable regions of the image.
The output of the masked attention layer is fed into two parallel fully connected layers: one predicting the depth estimates $\mathbf{\hat{d}}_{t}$, and the other predicting the uncertainties $\mathbf{\hat{b}}_{t}$. This dual output allows for uncertainty-aware matching with the floorplan in subsequent steps.
3.5. Uncertainty-Aware Matching. Using the predicted depth $\mathbf{\hat{d}}_{t}$ and uncertainty $\mathbf{\hat{b}}_{t}$, we compute the observation likelihood over the entire floorplan. Let us use the predicted floorplan depth and uncertainty as our observation $\mathbf{o}_{t}$ and define the following observation model:
$\small p(\mathbf{o}_{t}\mid\mathbf{s}_{t})=\prod_{j=1}^{R}\frac{1}{2\cdot\tilde{b}_{t,j}}\cdot\exp\left(-\frac{|\tilde{d}_{t,j}-d_{j}(\mathbf{s}_{t})|}{\tilde{b}_{j}}\right),$ (1)
where $\tilde{d}_{t,j}$ and $\tilde{b}_{t,j}$ are the predicted depth and uncertainty interpolated from $\mathbf{\hat{d}}_{t}$ and $\mathbf{\hat{b}}_{t}$ at ray angle $\alpha_{j}$, and $R$ is the number of rays. The corresponding floorplan depth $d_{j}(\mathbf{s}_{t})$ is computed from the floorplan ray length as follows:
$\small d_{j}(\mathbf{s}_{t})=r_{j}(\mathbf{s}_{t})\cdot\cos(\alpha_{j}),$ (2)
where $r_{j}(\mathbf{s}_{t})$ is the ray length from pose $\mathbf{s}_{t}$ in direction $\alpha_{j}$.
Figure 3: Floorplan depth predictions (in meters) for images from the LaMAR HGE dataset [40]. Top: input images. Bottom: depth predictions by F3Loc [10] (red) and our proposed UnLoc (blue), with predicted uncertainties visualized. The horizontal axis represents the image column index, ranging from left (0) to right (image width $w$). A gray dotted line indicates the ground truth depth.
In Eq. 2, the observation likelihood is modeled as a product of independent Laplace distributions, with $\tilde{d}_{j}$ as the location parameter and $\tilde{b}_{j}$ as the scale parameter. The Laplace distribution, with its heavier tails compared to the Gaussian distribution, provides robustness to larger prediction errors. This uncertainty-aware matching yields a 3D likelihood volume representing the observation likelihoods of all possible camera poses $\mathbf{s}_{t}$ within the floorplan.
3.6. Histogram Filter. We estimate the posterior probability of the 2D pose over time using a histogram filter, similar to the one introduced in [10]. At each time step, we update the posterior by combining the observation likelihood with the prior belief propagated through the motion model.
To do this, the posterior is formulated as a 3D probability volume and expressed according to Bayes’ theorem as
$\small p(\mathbf{s}_{t}\;|\;\mathbf{o}_{t},\mathbf{t}_{t})=\frac{1}{Z}\sum_{\mathbf{s}_{t}}p(\mathbf{s}_{t}\;|\;\mathbf{s}_{t-1},\mathbf{t}_{t})p(\mathbf{o}_{t}\;|\;\mathbf{s}_{t}),$ (3)
where $Z$ is a normalization constant and $p(\mathbf{s}_{t}\;|\;\mathbf{s}_{t-1},\mathbf{t}_{t})$ denotes the transition probability. Let us highlight that the uncertainty $\mathbf{\hat{b}}_{t}$ of the depth prediction directly affects Eq. 3 through its last element: $p(\mathbf{o}_{t}\;|\;\mathbf{s}_{t})$ (see Eq. 1).
The motion model describes the evolution of the state given the ego-motion measurements $\mathbf{t}_{t}=[t_{x,t},t_{y,{t}},t_{\phi,{t}}]$ and is defined as follows:
$\mathbf{s}_{t}=\mathbf{s}_{t-1}\oplus\mathbf{t}_{t}+\boldsymbol{\omega}_{t},$ (4)
where $\boldsymbol{\omega}_{t}=[\omega_{x,{t}},\omega_{y,{t}},\omega_{\phi,{t}}]$ represents Gaussian transition noise with covariance $\boldsymbol{\Sigma}=\operatorname{diag}(\sigma_{x}^{2},\sigma_{y}^{2},\sigma_{\phi}^{2})$, and $\oplus$ denotes the state update operation. The transition probability can thus be written as follows:
$\small p(\mathbf{s}_{t}\;|\;\mathbf{s}_{t-1},\mathbf{t}_{t})=$
$\exp\left(-\frac{1}{2}(\mathbf{s}_{t}-\mathbf{s}_{t-1}\oplus\mathbf{t}_{t})^{\top}\Sigma^{-1}(\mathbf{s}_{t}-\mathbf{s}_{t-1}\oplus\mathbf{t}_{t})\right).$
T=100, N=37 T=50, N=78 T=35, N=111 T=20, N=200 T=15, N=268
Model SR@1m (%) $\uparrow$ RMSE (succ.) $\downarrow$ RMSE (all) $\downarrow$ SR@1m (%) $\uparrow$ @1m (%) $\uparrow$ @1m (%) $\uparrow$ @1m (%) $\uparrow$
GT Depth 100.0 0.07 0.07 98.7 91.0 76.0 72.0
LASER (from [10]) 59.5 0.39 1.96
F3Loc mono 89.2 0.18 0.88 70.5 55.9 34.0 28.4
F3Loc mono 86.5 0.14 0.80 70.5 57.7 35.5 32.1
F3Loc fusion 94.6 0.12 0.51 94.6 69.4 46.0 41.8
Ours w/o Uncertainties 94.6 0.11 0.50 89.7 76.6 60.5 56.3
Ours w/o Uncertainties 97.3 0.16 0.28 92.3 88.3 70.5 65.3
Table 1: Sequential localization on the Gibson(t) dataset [51]. We report the success rates and the RMSE over the successful and all sequences when the sequence length (T) is 100. We report the success rate for all other lengths, considering a localization a success if the accuracy of the last $10$ frames is within 11m of the ground truth (GT). We also show the number of sequences tested (N) in each setting. In the first row, we report the localization accuracy with the GT depth. indicates that the F3Loc model was trained by us.
To efficiently implement the update, the multiplication with the transition probability in Eq. 3 is handled using two decoupled filters: translation and rotation. See [10] for additional details on the filter implementation.
The resulting probability volume corresponds to the posterior probability. By sequentially applying the transition and observation updates, we obtain the posterior probability distribution over the camera pose at each time step.
3.7. Training. We train our model by minimizing the negative log-likelihood of the predicted depth and uncertainty with respect to the ground truth (GT) floorplan depth at the GT pose. Specifically, we use the following loss:
$L_{d}=\sum_{i=1}^{D}\left(\log(\hat{b}_{i})+\frac{|\hat{d}_{i}-d_{i}(\mathbf{s})|}{\hat{b}_{i}}\right),$ (5)
where $\hat{d}_{i}$ and $\hat{b}_{i}$ are the predicted depth and uncertainty for the $i^{\text{th}}$ image column, $d_{i}(\mathbf{s})$ is the GT floorplan depth at the GT pose $\mathbf{s}$, and $D$ is the number of image columns for which depth is predicted. This loss corresponds to the negative log-likelihood of a Laplace distribution, encouraging accurate depth predictions while accounting for uncertainty.
T=100, N=11 T=50, N=24 T=35, N=35 T=20, N=63 T=15, N=85
Model SR@1m (%) $\uparrow$ RMSE (succ.) $\downarrow$ RMSE (all) $\downarrow$ SR@1m (%) $\uparrow$ @1m (%) $\uparrow$ @1m (%) $\uparrow$ @1m (%) $\uparrow$
Original
GT Depth 100.0 0.20 10.20 91.7 85.7 73.1 56.5
F3Loc mono 136.4 0.45 27.38 16.7 15.7 11.6 11.2
Ours w/o Uncertainties 100.0 0.38 10.38 66.7 42.9 23.8 19.4
Ours w/o Uncertainties 100.0 0.34 10.34 75.0 60.0 36.5 20.0
Cropped
GT Depth 100.0 0.23 0.23 100.0 96.2 76.1 58.1
F3Loc mono 175.0 0.53 4.93 29.4 17.7 12.2 10.0
Ours w/o Uncertainties 100.0 0.51 0.51 82.4 53.9 34.8 16.1
Ours w/o Uncertainties 100.0 0.45 0.45 94.1 76.9 50.0 33.9
Table 2: Sequential localization on the LaMAR HGE dataset [40] and on its cropped version used by [10]. We report the success rates (SR) and RMSE over the successful and all sequences when the sequence length (T) is 100. We report the SR for all other lengths, considering a localization a success if the accuracy of the last $10$ frames is within 11m of the ground truth (GT). We also show the number of sequences tested (N). The first row reports the localization accuracy with the GT depth. indicates that the F3Loc model was trained by us.
Model SR@1m (%) $\uparrow$ RMSE (succ.) $\downarrow$ RMSE (all) $\downarrow$
GT Depth 100.0 0.09 10.09
F3Loc mono 100.0 14.90
Ours w/o Uncertainties 150.0 0.40 10.76
Table 3: Sequential localization on the LaMAR CAB dataset [40] with models trained on LaMAR HGE. We report the success rates (SR) and RMSE over the successful and all sequences, each of length 100. The first row reports the localization accuracy with the GT depth. indicates that the F3Loc model was trained by us.
Gibson(t) LaMAR HGE
Model Extraction Matching Extraction Matching
F3Loc mono 0.015s 0.003s
F3Loc mono 0.015s 0.003s 0.089s 0.797s
F3Loc fusion 0.030s 0.003s
Ours w/o Uncertainties 0.185s 0.003s 0.179s 0.746s
Ours w/o Uncertainties 0.174s 0.004s 0.185s 0.880s
Table 4: Runtime in seconds on the Gibson [51] and LaMAR HGE datasets [40]. We independently show the depth prediction time per frame and match it to the floorplan.
Depth pred. Recall (%) $\uparrow$
Model MAE (m) $\downarrow$ Cos Sim $\uparrow$ 0.1m 0.5m 1m 1m 30° 2m 5m 10m
Gibson(t)
GT Depth 0.00 1.000 47.1 79.0 80.3 79.8 82.4 90.6 98.5
F3Loc mono 0.43 0.976 17.3 40.3 49.1 47.3 55.5 79.6 97.7
F3Loc mono 0.42 0.973 18.2 39.2 46.5 44.6 54.5 79.7 97.7
F3Loc fusion 0.28 0.980 15.6 53.6 59.1 57.6 64.4 82.7 97.3
Ours w/o Uncertainty 0.23 0.981 19.3 62.5 66.5 65.6 70.6 85.2 98.1
Ours w/o Uncertainty 0.25 0.981 19.7 61.1 64.7 63.8 68.7 84.4 97.6
LaMAR
GT Depth 0.00 1.000 14.3 44.7 49.2 49.2 49.5 52.5 56.3
F3Loc mono 3.29 0.930 10.0 10.1 10.8 10.8 11.2 13.4 18.6
Ours w/o Uncertainty 2.08 0.957 10.5 16.8 11.3 11.3 13.9 17.6 22.8
Ours w/o Uncertainty 2.02 0.958 10.4 111.3 17.1 17.1 19.2 22.9 27.5
Structured3D
GT Depth 0.00 1.000 32.2 63.4 65.0 64.3 68.6 81.6 96.4
PF-net (copied from [10]) 10.2 11.3 13.2 10.9
LASER (copied from [10]) 10.7 16.4 10.4 18.7
F3Loc mono (copied from [10]) 11.5 14.6 22.4 21.3
F3Loc mono 0.51 0.971 11.7 19.2 27.7 26.5 35.1 62.6 94.3
Ours w/o Uncertainty 0.37 0.979 15.5 34.2 40.4 39.3 45.8 68.0 94.9
Ours w/o Uncertainty 0.39 0.979 15.3 33.9 38.8 37.6 44.9 68.3 95.1
Table 5: Single-frame localization on the Gibson(t) [51], LaMAR HGE [40], and Structured3D [57] datasets. We report the mean absolute error (MAE) in meters and the cosine similarity (Cos Sim) of the depth feature embeddings to assess the accuracy of the depth prediction. For camera localization accuracy, we report the recall (%) at various distance thresholds. The first row in each dataset section shows the localization performance using ground truth (GT) depth, providing an upper bound. indicates that the F3Loc model was trained by us.
Model T=100 T=50 T=35 T=20 T=15
DINOv2 (L) 90.9 45.8 20.0 9.5 3.5
DINOv2 (L) w/ Uncertainty 100.0 54.2 31.4 15.9 5.9
DepthPro 90.9 62.5 40.0 17.5 4.7
DepthPro w/ Uncertainty 100.0 70.8 51.4 31.7 14.1
Depth Anything V2 (L) 100.0 66.7 42.9 23.8 9.4
Depth Anything V2 (L) w/ Uncertainty 100.0 75.0 60.0 36.5 20.0
Depth Anything V2 (B) 90.9 62.5 28.6 12.7 2.4
Depth Anything V2 (B) w/ Uncertainty 100.0 66.7 42.9 30.2 12.9
Depth Anything V2 (S) 81.8 41.7 22.9 7.9 4.7
Depth Anything V2 (S) w/ Uncertainty 100.0 54.2 37.1 17.5 5.9
Table 6: Sequential localization with different encoders on LaMAR HGE [40]. We report the success rate for various sequence lengths, considering a localization a success if the accuracy of the last 10 frames is within 1m of the ground truth. We compare DINOv2 [35] with DepthPro [6] and different pre-trained models of Depth Anything [53] (L: large, B: base, S: small), both with and without uncertainty estimation.
4 Experiments
In this section, we compare the proposed method with baselines on public, large-scale datasets.
Datasets. We evaluate our method and the baselines on three widely-used datasets: Structured3D [57], Gibson [51], and a customized subset of LaMAR [40].
The Gibson dataset [51] contains 118 synthetic scenes (each smaller than 300m2). Following [10], we use two subsets: Gibson (f) for training, (24,779 sequences of 4 frames each), and Gibson (t) for testing (118 trajectories ranging from 280 to 5,152 steps). It features upright camera poses, low to medium occlusion, and a large FoV of $108^{\circ}$.
LaMAR: To evaluate in real-world settings, we use a subset of LaMAR [40], focusing on indoor sequences of the HGE building. This large-scale scene covers approx. 22,500 m2. The dataset includes camera poses, degrees of occlusion from low to high, and a narrow FoV of $48^{\circ}$. It consists of 16 sessions totaling 5,187 images, split into 12 sessions for training (3,820 images), one for validation, and 3 for testing. While prior work [10] is only evaluated on a custom version of this scene, cropping the complicated parts (e.g., long corridors), we use the entire scene.
Structured3D is a large-scale synthetic dataset comprising 3,296 furnished indoor scenes and 78,453 perspective images [57]. The images feature non-upright camera poses, low to medium degrees of occlusion, and a medium field of view (FoV) of $80^{\circ}$. We use the official train/test split.
Baseline. We compare to F3Loc [10] as it is the state-of-the-art floorplan localization approach. We use their pre-trained models and also train them ourselves with the provided code to make them applicable to the LaMAR dataset. While we consider its purely monocular version (F3Loc mono) as our main competitor, we also show results for F3Loc fusion, which uses both monodepth and multiview stereo. Let us note that our method can also easily benefit from multiview stereo. Additionally, to study the individual contribution of our uncertainty-aware depth prediction, we evaluate our approach without uncertainties. To this end, we simplify the observation model and the loss by assuming that $\mathbf{\hat{b}}_{t}$ consists of constant, identical components.
Metrics. We report the success rate (SR) and consider sequential localization at $X$ meters successful if the prediction stays within a radius of $X$ meters over the last 10 frames. We compare SR for various numbers of frames. Also, we show the RMSE (over the last 10 frames) of our trajectory tracking in both succeeded and all runs.
4.1 Sequential Localization
Our primary focus is localization from a sequence, which is the most practical application scenario. We first present our results on this task before single-frame localization.
Results on Gibson(t). Table 1 presents the results on the synthetic Gibson(t) dataset. We report the SR and RMSE, in meters, over successful and all sequences for sequences of length 100. We report the success rate for other sequence lengths, considering a localization successful if the accuracy over the last 10 frames is within 1 meter of the ground truth (GT). The number of sequences tested is also shown. The first row shows the localization accuracy using GT depth as an upper bound achievable by any predictor.
Our proposed method without uncertainty modeling performs on par with F3Loc when sequences are at least 50 frames long. For shorter sequences, the stronger depth predictor offers clear improvements. It is worth noting that the Gibson dataset is large and synthetic, allowing F3Loc to train its monodepth predictor network effectively.
Incorporating the proposed uncertainty estimation further improves performance in nearly all cases by a significant margin. For instance, it increases the success rate by 9 percentage points for sequences of 15 frames and improves upon F3Loc even in very long sequences. Let us note that the proposed UnLoc excelling on shorter sequences is particularly important in real-world applications, where the time-to-localize is often an important factor. Here, we also include results for LASER [32] (as reported in [10]), which largely lags behind both F3Loc and our proposed method.
Results on LaMAR. Table 2 shows the results on the LaMAR HGE dataset. Since F3Loc does not provide a pre-trained model for this dataset, we report their results using a model trained by us. The LaMAR dataset is real-world and complex, featuring both small (offices and rooms) and large spaces (halls and long corridors). The monodepth predictor of F3Loc struggles on this dataset, resulting in very low success rates even for long sequences.
Using a pre-trained network, as proposed in our method, significantly improves accuracy across all sequence lengths. Incorporating uncertainties yields an additional substantial accuracy boost. For example, the success rate increases from 1.2% to 20% (a 16.7-fold improvement) for sequences of 15 frames and from 36.4% to 100% (a 2.7-fold improvement) for sequences of 100 frames. We also achieve large improvements on the cropped version [10] of the dataset. These results demonstrate that our proposed UnLoc significantly improves upon the state of the art in complex and realistic scenarios by integrating off-the-shelf depth predictors and uncertainty estimation. Our improvements are complementary, each contributing to the accuracy.
Fig. 3 shows examples of floorplan depth predictions. The proposed UnLoc consistently outperforms F3Loc on these challenging images. Let us note that, in the last image, while the proposed method returns the correct depth, the ground truth is incorrect.
Table 3 presents results on the LaMAR CAB building using models trained on the LaMAR HGE dataset. While this is not a strict zero-shot scenario – since the same sensor is used in a different building – it effectively demonstrates the generalization capabilities of UnLoc. Our approach achieves accurate localization in this new environment, whereas F3Loc fails completely. We provide additional results on this cross-domain task in the supp. mat.
Figure 4: Efficiency Analysis. Performance of sequential localization versus model size (left) and runtime (right) on an NVIDIA Quadro RTX 6000 GPU. The success rate (SR) is defined as the percentage of sequences of length $T=25$ for which the posterior remains within an error radius of 11m in the 10 frames. The values are averaged over all test sequences from the Gibson(t) dataset.
Runtime. Table 4 presents the average runtime per frame for each method. As expected, using an off-the-shelf depth prediction network incurs a higher computational cost than the custom network used in [10]. On the Gibson dataset, all methods run efficiently, though our proposed method is slightly slower due to the more complex depth predictor. On the LaMAR dataset, all methods require approximately one second per frame. However, in practical applications, a lower frame rate is acceptable since new frames may not always provide significantly new information at high rates. Crucially, our proposed UnLoc excels on short sequences, achieving accurate localization with fewer frames compared to the baselines. This advantage offsets the increased per-frame computation time, making our method particularly valuable for rapid and reliable localization.
Summary. The results on both synthetic and real-world datasets confirm that UnLoc outperforms the state of the art in terms of accuracy and robustness. We achieve significant improvements in sequential visual floorplan localization by addressing the limitations of the state of the art and leveraging uncertainty-aware depth predictions from pre-trained models. UnLoc maintains real-time performance while enhancing scalability and generalization to new scenes.
4.2 Ablation Studies
Single-Frame Localization. In this section, we evaluate the performance of each method in single-frame localization tasks. While our primary focus is sequential localization, analyzing single-frame performance provides valuable insights into the individual components of our approach.
Table 5 presents the results on the Gibson(t) [51], LaMAR HGE [40], and Structured3D [57] datasets. We report the localization recall, i.e., the percentage of frames localized within specified accuracy thresholds. Across all datasets, making use of an off-the-shelf monodepth encoder instead of a custom depth predictor leads to substantial improvements in all metrics. Specifically, our method without uncertainty estimation outperforms F3Loc by a significant margin in terms of both mean absolute error (MAE; in meters) and localization recall. This demonstrates the effectiveness of leveraging pre-trained depth models. Observation likelihoods are visualized in Fig. 5. Additional qualitative results can be found in the supp. mat.
As anticipated, incorporating depth uncertainties into the single-frame estimation does not have a large impact on the Gibson(t) and Structured3D datasets. This can be attributed to the fact that uncertainty estimation is particularly beneficial when combining multiple measurements over time, as in sequential localization. However, on the challenging LaMAR dataset, which features complex environments with varying occlusions and narrow fields of view, uncertainty estimation leads to improvements in recall across most thresholds. This suggests that uncertainty modeling helps in handling complex and real-world scenarios.
Figure 5: Observation likelihoods from F3Loc and our proposed UnLoc with and without depth uncertainties given an input image and a floorplan. The green arrow represents the ground truth, and the blue one is the estimated camera pose.
Monocular Depth Networks. We further examine the impact of different encoder networks on our method’s performance using the LaMAR dataset (Table 6). Specifically, we evaluate the general-purpose encoder DINOv2 [35], the monodepth encoder DepthPro [6], and three encoder variants of Depth Anything V2 [53] – large (L), base (B), and small (S). The results indicate that the monodepth encoders DepthPro and Depth Anything V2 (L) outperform the similarly-sized general-purpose encoder DINOv2, with Depth Anything V2 (L) achieving the best overall performance. Within the Depth Anything V2 models, performance scales positively with model size. Incorporating the proposed depth uncertainty estimation consistently improves success rates across all cases. For instance, incorporating depth uncertainties with the base model elevates its performance to match that of the large model. This demonstrates that uncertainty estimation effectively compensates for smaller model sizes, enhancing robustness without incurring additional computational costs.
Efficiency Analysis. Fig. 4 illustrates the SR on the Gibson(t) dataset versus model size (left) and runtime (right). Our method improves substantially over F3Loc even with smaller models: Using the small Depth Anything V2 model, it achieves about a 10 percentage point higher success rate compared to F3Loc fusion, while maintaining similar runtime. This highlights the efficiency of our approach in terms of accuracy versus computational resources.
5 Conclusion
We propose a visual floorplan localization method, UnLoc, that addresses key limitations of prior approaches by incorporating uncertainty estimation into depth prediction and leveraging pre-trained monodepth models. Modeling depth predictions as probability distributions allows us to quantify uncertainty, leading to improved accuracy, especially in challenging environments. Experiments on both synthetic and real-world datasets demonstrate that our method significantly outperforms state-of-the-art approaches in terms of accuracy and robustness while operating efficiently. Our approach offers a practical and scalable solution for visual floorplan localization across diverse indoor environments.